OpenStudy (adajiamcneal):

What is the sum of the interior angle measures of a 37-gon Explain

OpenStudy (adajiamcneal):

@Here_to_Help15

OpenStudy (here_to_help15):

Is this a multiple choice or essay answers?

OpenStudy (adajiamcneal):

essay

OpenStudy (here_to_help15):

Well to begin with (2n-4)*90 and in this problem n=37 so (2n-4)90 =(2*37-4)90=(74-4)90=70*90=6300

OpenStudy (adajiamcneal):

thanks :)